# Random-Set Neural Networks (RS-NN)

### v2

@@ 3 Random-Set Neural Network
3.1 Architecture
Representation. A classifier $e$ (e.g., a neural network) is a mapping from an input space $X$ to a target space $\mathcal{C}$ (the set of classes), i.e., $e:X\rightarrow\mathcal{C}$. In our set-valued classification setting, on the other hand, $e$ is a mapping from $X$ to the set of all subsets of $\mathcal{C}$, the powerset $\mathbb{P}(\mathcal{C})$, namely: $e:X\rightarrow\mathbb{P}(\mathcal{C})$.
@@ 3.1 Architecture
As shown in Fig. 5, RS-NN predicts for each input data point a belief function, rather than a vector of softmax probabilities as in a traditional CNN. For $N$ classes, a ‘vanilla’ RS-NN would have $2^{N}$ outputs (as $2^{N}$ is the cardinality of $\mathbb{P}(\mathcal{C})$), each being the belief value (2) of the focal set of classes $A\in\mathbb{P}(\mathcal{C})$ corresponding to that output neuron. Given a training datapoint with attached a true class, its ground truth is encoded by the vector $\mathbf{bel}=\{Bel(A),A\in\mathbb{P}(\mathcal{C})\}$ of belief values for each focal set of classes $A\in\mathbb{P}(\mathcal{C})$, where $Bel(A)$ is set to 1 iff the true class is contained in the subset $A$, 0 otherwise. [^1] Note that the belief encoding of ground-truth is not related to label smoothing. It maps ground-truth labels from the original class space to a set space without adding noise, thus preserving label ”hardness”. This corresponds to full certainty that the element belongs to that set and there is complete confidence in this proposition.
@@
Figure 5: RS-NN model architecture. Given a collection $\mathcal{C}$ of $N$ classes, the network selects the top $K$ relevant (focal) sets of classes $A_{1},\ldots,A_{K}$ from the powerset $\mathbb{P}(\mathcal{C})$, encodes the ground truth training classes as belief functions $Bel$, and uses them to predict a belief function $\mathbf{\hat{bel}}$ for each input test data point, using the loss (7). Mass values $\hat{m}$ and pignistic probability estimates $BetP$ are computed from the predicted belief function. Uncertainty is estimated as described in Sec. 3.3
@@
Budgeting. To overcome the exponential complexity of using $2^{N}$ sets of classes (especially for large $N$), a fixed budget of $K$ relevant non-singleton (of cardinality $>1$) focal sets are used. These focal sets are obtained by clustering the original classes $\mathcal{C}$ and selecting the top $K$ focal sets of classes with the highest overlap ratio, computed as the intersection over union for each subset $A$ in $\mathbb{P}(\mathcal{C})$: $overlap(A)={\cap_{c\in A}A^{c}}/{\cup_{c\in A}A^{c}},\>A_{1},\ldots,A_{K}\in\mathbb{P}(\mathcal{C})$. The clustering is performed on feature vectors of images of each class generated by a standard CNN trained on the original classes $\mathcal{C}$. The feature vectors are further reduced to 3 dimensions using t-SNE (t-Distributed Stochastic Neighbor Embedding) JMLR:v9:vandermaaten08a before applying a Gaussian Mixture Model (GMM) to them. Ellipsoids spruyt2014draw , covering $95\%$ of data, are generated using eigenvectors and eigenvalues of the covariance matrix $\Sigma_{c}$ and the mean vector $\mu_{c}$, $\forall c\in\mathcal{C},P_{c}\sim\mathcal{N}(x_{c};\mu_{c},\Sigma_{c})$ obtained from the GMM to calculate the overlaps. To avoid computing a degree of overlap for all $2^{N}$ subsets, the algorithm is early stopped when increasing the cardinality does not alter the list of most-overlapping sets of classes.
The $K$ non-singleton focal sets so obtained, along with the $N$ original (singleton) classes, form our network outputs $\mathcal{O}=\mathcal{C}\cup\{A_{1},\ldots,A_{K}\}$. E.g., in a $100$-class scenario, the powerset contains $2^{100}$ subsets ($10^{30}$ possibilities). Setting a budget of $K=200$, for instance, results in $300100+K=$ 300 outputs, a far more manageable number.
The budgeting step is a one-time procedure, applied to any given dataset before training. It requires just 2 minutes for CIFAR-10, 7 minutes for CIFAR-100, and 2 hours for ImageNet ($\approx$ 1.1M images, 1000 classes). In return, it greatly helps overcome the computational challenge of dealing with a large number of classes. Budgeting can be performed on features obtained from pre-trained classifiers.
@@ 3.2 Loss function
A Random-Set prediction problem is mathematically similar to the multi-label classification problem, for in both cases the ground truth vector contains several 1s. In the former case, these correspond to sets all containing the true class; in the latter, to all the class labels attached to same data point.
Despite the different semantics, we can thus adopt as loss Binary Cross-Entropy (BCE) (5) with sigmoid activation, to drive the prediction of a belief value for each focal set in the identified budget.
$\begin{array}[]{lll}\mathcal{L}_{BCE}=\displaystyle-\sum_{i=1}^{b_{size}}\frac{1}{|\mathcal{O}|}\sum_{A\in\mathcal{O}}\Big{[}Bel_{i}(A)\log(\hat{Bel}_{i}(A))+(1-Bel_{i}(A))\log(1-\hat{Bel}_{i}(A))\Big{]},\end{array}$ (5)
where $i$ is the index of the training point within a batch of cardinality $b_{size}$, $A$ is a focal set of classes in the budget $\mathcal{O}$, $Bel_{i}(A)$ is the $A$-th component of the vector $\mathbf{bel}_{i}$ encoding the ground truth belief values for the $i$-th training point, and $\hat{Bel}_{i}(A)$ is the corresponding belief value in the predicted vector $\hat{\mathbf{bel}}_{i}$ for the same training point. Both $\mathbf{bel}_{i}$ and $\hat{\mathbf{bel}}_{i}$ are vectors of cardinality $|\mathcal{O}|$ for all $i$.
To ensure that the predicted scores amount to a valid belief function, we incorporate a mass regularization term $M_{r}$ in the loss function that encourages non-negativity (1) of the (predicted) mass values $\hat{m}(A)$, $A\in\mathcal{O}$, which can be obtained from the predicted belief function $\hat{\mathbf{bel}}$ via the Moebius inversion formula (2). For it to be valid, the sum of the masses of the predicted belief function must be equal to 1 (1), therefore, we also add a mass sum term $M_{s}$, namely
$\begin{array}[]{l}\displaystyle M_{s}=\max\left(0,\frac{1}{b_{size}}\sum_{i=1}^{b_{size}}\sum_{A\in\mathcal{O}}\hat{m}_{i}(A)-1\right),\hskip 5.0pt\displaystyle M_{r}=\frac{1}{b_{size}}\sum_{i=1}^{b_{size}}\sum_{A\in\mathcal{O}}\max(0,-\hat{m}_{i}(A)).\end{array}$ (6)
All the components $\mathcal{L}_{BCE}$, $M_{r}$ and $M_{s}$ are computed during batch training. Two hyperparameters, $\alpha$ and $\beta$, control the relative importance of the two mass terms, yielding as the total loss for RS-NN:
@@
The regularization terms aim to penalize deviations from valid belief functions. When the values of $\alpha$ and $\beta$ are too small, the regularization terms may not be sufficient to ensure that predictions are valid belief functions. This occurs because we do not enforce the model to predict valid belief functions during training; rather, we merely encourage it – just like, e.g., training time regularisation in neurosymbolic learning encourages predictions to be commonsense giunchiglia2023road . In such cases,  [^2] At any rate, improper belief functions are employed in the literature, e.g. for conditioning cuzzolin2020geometry post-training, we fix all negative masses (obtained from predicted belief functions) to zero and introduce a final subset, often referred to as the ‘universal set’ or ‘full set,’ containing all the classes in the powerset, to the final budget $\mathcal{O}$. This subset is assigned all the remaining mass, ensuring that the sum of masses across all focal sets in $\mathcal{O}$ equals 1. This adjustment compensates for the fact that, during clustering, only a few focal sets of classes $\{A_{1},\ldots,A_{K}\}$ from the power set $\mathbb{P}(\mathcal{C})$ are selected, rather than considering any arbitrary set of classes.
@@ 3.3 Accuracy and uncertainty estimation
Pignistic prediction. The pignistic probability (3) is the central prediction associated with a belief function (seen as a credal set): standard performance metrics such as accuracy can then be calculated after extracting the most likely class according to the pignistic prediction (e.g. in §D.5.3).
@@
$H_{RS}=-\sum_{c\in\mathcal{C}}{BetP}(c)\log{BetP}(c).$ (8)
@@
Size of the credal prediction. A sensible measure of the epistemic uncertainty attached to a random-set prediction $\mathbf{\hat{bel}}$ is the size of the corresponding credal set (4). There are several ways of measuring the size of a convex polytope like a credal set sale2023volume . Given the upper and lower bounds to the probability assigned to each class $c$ by distributions within the predicted credal set $\hat{Cre}$,
$\overline{P}(c)=\max_{P\in\hat{Cre}}P(c),\quad\underline{P}(c)=\min_{P\in\hat{Cre}}P(c),$ (9)
we propose a simple way to measure the size of the credal set as the difference between the lower and upper bounds (9) associated with the most likely class according to the pignistic prediction. For each class $c$, not just $\hat{c}$, the predicted pignistic estimate $BetP(c)$ falls within the interval $[\underline{P}(c),\overline{P}(c)]$, whose width $\overline{P}(c)-\underline{P}(c)$ indicates the epistemic uncertainty associated with the prediction (§A.4).
@@ 4.1 Implementation
Experiments, datasets and baselines. Firstly, we assess the test accuracy (%) and inference time (ms) on multi-class image classification datasets, including MNIST LeCun2005TheMD , CIFAR-10 cifar10 , Intel Image intelimage , CIFAR-100 krizhevsky2009learning , and ImageNet deng2009imagenet . Tab. 2 provides a comparison of test accuracies and inference times of RS-NN against state-of-the-art Bayesian methods, R-BNN https://doi.org/10.48550/arxiv.1312.6114 , LB-BNN hobbhahn2022fast , Ensemble classifier ENN osband2024epistemic and standard CNN as detailed in Sec. 4.2.
Our second set of experiments (Sec. 4.3) concerns out-of-distribution (OoD) detection. Tab. 1 shows our results on OoD detection metrics AUROC (Area Under Receiver Operating Characteristic curve) and AUPRC (Area Under Precision-Recall curve) for RS-NN, LB-BNN, ENN and standard CNN. We provide quantitative assessment on several iD vs OoD datasets: CIFAR-10 vs SVHN netzer2011reading /Intel-Image intelimage , MNIST vs F-MNIST xiao2017fashion /K-MNIST clanuwat2018deep , and ImageNet vs ImageNet-O hendrycks2021natural . We also report the Expected Calibration Error (ECE) for RS-NN, LB-BNN, ENN and standard CNN on datasets CIFAR-10, MNIST and ImageNet (see Tab. 1).
In the third set of experiments (Sec. 4.4), we test the uncertainty estimation capabilities of RS-NN using both the entropy of the pignistic prediction and the size of the predicted credal set. Tab. 1 presents the predicted pignistic entropy of RS-NN alongside entropies of LB-BNN, ENN, and CNN on in-distribution (iD) datasets, namely CIFAR-10, MNIST, ImageNet, and out-of-distribution (OoD) datasets CIFAR-10 vs SVHN/Intel Image, MNIST vs F-MNIST/K-MNIST, and ImageNet vs ImageNet-O. Tab. 3 shows credal set widths for RS-NN predictions across the same datasets.
In our final set of experiments, we explore the adaptability of RS-NN (Sec. 4.5) to large-scale architectures, employing it on models such as WideResNet-28-10, VGG16, Inception V3, EfficientNetB2 and ViT-Base. Further, to underscore the model’s ability to leverage transfer learning, we conduct training and testing on a pre-trained ResNet50 model initialised with ImageNet weights (Tab. 4).
Additional experiments include providing statistical guarantees on RS-NN using non-conformity scores (§A), evaluating the robustness of RS-NN to adversarial attacks (§D.2), noisy and rotated in-distribution samples in Appendix §D.1, how RS-NN circumvents the overconfidence problem in CNNs is given in §D.3. Results showing how credal set width is less correlated with confidence scores than entropy are discussed in in §D.5.1. Ablation studies on $\alpha$ and $\beta$ are shown in §D.5.5 and number of non-singleton focal sets ($K$) in §D.5.6. The ablation study demonstrates that the best accuracy is obtained at small values (0 to 1e-2) of $\alpha$ and $\beta$, and $K=20$ for CIFAR-10 dataset.
@@
Baselines and backbones. All models, including RS-NN, R-BNN, LB-BNN, ENN and standard CNN, are trained on ResNet50 (on 3 NVIDIA A100 80GB GPUs) with a learning rate scheduler initialized at 1e-3 with 0.1 decrease at epochs 80, 120, 160, and 180. Standard data augmentation NIPS2012_c399862d including random horizontal/vertical shifts with a magnitude of 0.1 and horizontal flips are applied to all models. ResNet50, excluding the top classification layer, serves as the common architecture. ResNet was originally designed for classification on the ImageNet dataset (1000 classes). To accommodate a reduced number of classes in smaller datasets (e.g., CIFAR-10/MNIST with 10 classes, Intel Image with 7 classes, and CIFAR-100 with 100 classes), two additional dense layers (1024 and 512 neurons, ReLU activation) are added to ResNet50. Similar techniques are commonly applied in deep learning to adapt model architectures to different datasets He_2016_CVPR ; zagoruyko2016wide . The output layer of RS-NN on ResNet50 has the same number of units as the number of (selected) focal sets $|\mathcal{O}|$, and uses sigmoid activation since the ground truth encoding resembles a multi-label classification problem (see Sec. 3.2). For all other models, the final output layer simply consists of a softmax activation for multi-class classification.
@@
Training the RS-NN. We first train a CNN using Adam kingma2014adam optimizer on categorical cross-entropy loss over 100 epochs with learning rate scheduler and batch size of 128. The feature vector $f$ is extracted from the penultimate layer of this trained model for clustering. Alternatively, features can be obtained directly from a pre-trained classifier. In the clustering phase, we use 50 CPU cores (all available), setting a budget $K$ of 20 focal sets for CIFAR-10/ MNIST/ Intel Image, 200 for CIFAR-100 and 3000 for ImageNet. RS-NN is trained from scratch on ground-truth belief encoding of sets using the $\mathcal{L}_{RS}$ loss function (7) over 200 epochs, with a batch size ($b_{size}$) of 128. The data is split into 40000:10000:10000 samples for training, testing, and validation respectively for CIFAR-10 and CIFAR-100, 50000:10000:10000 samples for MNIST, 13934:3000:100 samples for Intel Image, 1172498:50000:108669 for ImageNet. For all OoD datasets, we use 10,000 testing samples except for Intel Image (CIFAR-10 vs Intel Image), we use 3000 samples. All datasets are resized to $224\times 224$.
@@
RS-NN $\mathbf{99.71}$ Pignistic entropy $0.010\pm 0.111$ $\mathbf{0.0029}$ $\mathbf{93.89}$ $\mathbf{93.98}$ $\mathbf{0.530}\pm\mathbf{0.770}$ $\mathbf{96.75}$ $\mathbf{96.58}$ $\mathbf{0.740}\pm\mathbf{0.917}$
@@
ENN $99.07$ Mean Entropy $0.022\pm 0.127$ $0.0039$ $81.79$ $82.92$ $0.313\pm 0.464$ $95.94$ $95.45$ $0.703\pm 0.672$

### v3

@@ 3 Random-Set Neural Network
3.1 Approach
Representation. As shown in Fig. 5 (b), RS-NN predicts for each input data point a belief function, rather than a vector of softmax probabilities as in a traditional CNN. For $N$ classes, a ‘vanilla’ RS-NN would have $2^{N}$ outputs (as $2^{N}$ is the cardinality of $\mathbb{P}(\mathcal{C})$), each being the belief value (Eq. 2) of the focal set of classes $A\in\mathbb{P}(\mathcal{C})$ corresponding to that output neuron. Our architecture focuses primarily on the final layers (Fig. 5 (b), in grey), acting as a wrapper applicable on top of any baseline representation layers from existing models (Fig. 5 (b), in blue). This enables the integration of pre-trained networks while fine-tuning only the final decision layers. Hence, RS-NN is easily scalable to any model architecture, as demonstrated in Sec. 4.5, Tab. 4.
@@ 3.1 Approach
Given a training datapoint with a true class attached, its ground truth is encoded by the vector $\mathbf{bel}=\{Bel(A),A\in\mathbb{P}(\mathcal{C})\}$ of belief values for each focal set of classes $A\in\mathbb{P}(\mathcal{C})$. $Bel(A)$ is set to 1 iff the true class is contained in the subset $A$, 0 otherwise [^1] Note that the belief encoding of ground-truth is not related to label smoothing. It maps ground-truth labels from the original class space to a set space without adding noise, thus preserving label “hardness”.. This corresponds to full certainty that the element belongs to that set and there is complete confidence in this proposition (see §A.2 for an example).
@@
Figure 5: RS-NN model architecture. (a) Budgeting: Given a collection $\mathcal{C}$ of $N$ classes, the top $K$ relevant (focal) sets of classes $\{A_{1},\ldots,A_{K}\}$ are selected from the powerset $\mathbb{P}(\mathcal{C})$ (Algorithm in §C.1) and added to the singleton classes to form the budget $\mathcal{O}$. (b) Training and Inference: Ground truth classes are encoded as belief vectors $\mathbf{bel}$ and used to predict a belief function $\mathbf{\hat{bel}}$ for each training data point by minimising the loss $\mathcal{L}_{RS}$ (7), to train the output layers (in grey) producing $\mathbf{\hat{bel}}$. Mass values $\hat{m}$ and pignistic probability estimates $BetP$ are computed from the predicted belief function. Uncertainty is estimated as described in Sec. 3.3.
@@
Budgeting. To overcome the exponential complexity of using $2^{N}$ sets of classes (especially for large $N$), a fixed budget of $K$ relevant non-singleton (of cardinality $>1$) focal sets are used. These focal sets are obtained by clustering the original classes $\mathcal{C}$, fitting ellipses over them, and selecting the top $K$ focal sets of classes with the highest overlap ratio (Fig. 5 (a)). This is computed as the intersection over union for each subset $A$ in $\mathbb{P}(\mathcal{C})$: $overlap(A)={\cap_{c\in A}A^{c}}/{\cup_{c\in A}A^{c}},\>A_{1},\ldots,A_{K}\in\mathbb{P}(\mathcal{C})$. Clustering is performed on feature vectors of images generated by any feature extractor trained on the original classes $\mathcal{C}$. In our experiments, we used features from a trained standard ResNet50 model. These feature vectors are further reduced to 3 dimensions using t-SNE (t-Distributed Stochastic Neighbor Embedding) (van der Maaten & Hinton, 2008) before applying a Gaussian Mixture Model (GMM) to them. Ellipsoids (Spruyt, 2013), covering $95\%$ of data, are generated using eigenvectors and eigenvalues of the covariance matrix $\Sigma_{c}$ and the mean vector $\mu_{c}$, $\forall c\in\mathcal{C},P_{c}\sim\mathcal{N}(x_{c};\mu_{c},\Sigma_{c})$ obtained from the GMM to calculate the overlaps. To avoid computing a degree of overlap for all $2^{N}$ subsets, the algorithm is early stopped when increasing the cardinality does not alter the list of most-overlapping sets of classes. The $K$ non-singleton focal sets so obtained, along with the $N$ original (singleton) classes, form our network’s budget of outputs $\mathcal{O}=\mathcal{C}\cup\{A_{1},\ldots,A_{K}\}$. E.g., in a $100$-class scenario, the powerset contains $2^{100}$ subsets ($10^{30}$ possibilities). Setting a budget of $K=200$, for instance, results in $300100+K=$ 300 outputs, a far more manageable number.
As shown in our experiments, budgeting also helps the network converge without overfitting. While, in theory, a complete belief function model would be more powerful, in practice a larger number of focal sets impedes the learning process. As shown in Tab. 16, §E.7, RS-NN with a limited number of well-representative sets often performs better than RS-NN with a full power set of classes, and is more efficient at uncertainty estimation. The budgeting step is a one-time procedure, applied to any given dataset before training. It requires just 2 minutes for CIFAR-10, 7 minutes for CIFAR-100, and 60 minutes for ImageNet ($\approx$ 1.1M images, 1000 classes).
Further, our budgeting procedure is not specific to t-SNE, but can use any dimensionality reduction technique (Maćkiewicz & Ratajczak, 1993). For instance, Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2018) requires approximately 1 minute for the CIFAR-10 dataset, 2 minutes for CIFAR-100, and around 23 minutes for ImageNet to generate embeddings (see Tabs. 17, 18 in §E.8 for a t-SNE vs UMAP ablation study). This is including the overlap computation which only takes a few seconds and is efficiently parallelised across 150 CPU cores.
@@ 3.2 Loss function
A random-set prediction problem is mathematically similar to the multi-label classification problem, for in both cases the ground truth vector contains several 1s. In the former case, these correspond to sets all containing the true class; in the latter, to all the class labels attached to same data point. Despite the different semantics, we can thus adopt as loss Binary Cross-Entropy (BCE) (5) with sigmoid activation, to drive the prediction of a belief value for each focal set in the identified budget:
$\begin{array}[]{lll}\mathcal{L}_{BCE}=\displaystyle-\frac{1}{b_{size}}\sum_{i=1}^{b_{size}}\frac{1}{|\mathcal{O}|}\sum_{A\in\mathcal{O}}\Big{[}Bel_{i}(A)\log(\hat{Bel}_{i}(A))+(1-Bel_{i}(A))\log(1-\hat{Bel}_{i}(A))\Big{]}.\vspace{-5pt}\end{array}$ (5)
Here, $i$ is the index of the training point within a batch of cardinality $b_{size}$, $A$ is a focal set of classes in the budget $\mathcal{O}$, $Bel_{i}(A)$ is the $A$-th component of the vector $\mathbf{bel}_{i}$ encoding the ground truth belief values for the $i$-th training point, and $\hat{Bel}_{i}(A)$ is the corresponding belief value in the predicted vector $\hat{\mathbf{bel}}_{i}$ for the same training point. Both $\mathbf{bel}_{i}$ and $\hat{\mathbf{bel}}_{i}$ are vectors of cardinality $|\mathcal{O}|$ for all $i$.
A valid belief function satisfies Eq. 1 which states that mass values derived from belief functions should be non-negative and should sum up to 1 (Shafer, 1976). To ensure this, we incorporate a mass regularization term $M_{r}$ and a mass sum term $M_{s}$ in the loss function:
$\begin{array}[]{l}\displaystyle M_{r}=\frac{1}{b_{size}}\sum_{i=1}^{b_{size}}\sum_{A\in\mathcal{O}}\max(0,-\hat{m}_{i}(A)),\hskip 5.0pt\displaystyle M_{s}=\max\left(0,\frac{1}{b_{size}}\sum_{i=1}^{b_{size}}\sum_{A\in\mathcal{O}}\hat{m}_{i}(A)-1\right).\end{array}$ (6)
$M_{r}$ encourages non-negativity of the (predicted) mass values $\hat{m}(A)$, $A\in\mathcal{O}$. These mass values are obtained from the predicted belief function $\hat{\mathbf{bel}}$ via the Moebius inversion formula (Eq. 2). For it to be valid, the sum of the masses of the predicted belief function must be equal to 1 (Eq. 1), which is encouraged by the mass sum term $M_{s}$. All loss components $\mathcal{L}_{BCE}$, $M_{r}$ and $M_{s}$ are computed during batch training. Two hyperparameters, $\alpha$ and $\beta$, control the relative importance of the two mass terms, yielding as the total loss for RS-NN:
@@
The regularisation terms aim to penalise deviations from valid belief functions, in line with, e.g., the way training time regularisation in neurosymbolic learning encourages predictions to be commonsense (Giunchiglia et al., 2023). In general, soft constraints (Márquez-Neila et al., 2017) have been shown to not be inferior to hard ones. Still, when $\alpha$ and $\beta$ are too small, this may not be sufficient to ensure that predictions are valid belief functions. In such cases, [^2] At any rate, improper belief functions are normally used in the literature (Denoeux, 2021), e.g. for conditioning (Cuzzolin, 2020). post-training, we set any negative masses to zero and add the ‘universal’ set of all classes to the final budget. This subset is assigned all the remaining mass, ensuring that the sum of masses across all focal sets in $\mathcal{O}$ equals 1. This approach mimics classical approximation schemes (e.g. Cuzzolin (2020), Part III).
@@ 3.3 Accuracy and uncertainty estimation
Pignistic prediction. The pignistic probability (Eq. 3) is the central prediction associated with a belief function (seen as a credal set): standard performance metrics such as accuracy can then be calculated after extracting the most likely class according to the pignistic prediction (e.g. in §E.5.1). Notably, the RS-NN architecture and training mechanism are designed to facilitate set-based learning by incorporating class set information during training. When pignistic probabilities are computed during inference, they reflect the learning derived from the masses and beliefs of various subsets, making it more reliable than traditional softmax probabilities, as shown in §E.2.
@@
$H_{RS}=-\sum_{c\in\mathcal{C}}{BetP}(c)\log{BetP}(c).\vspace{-8pt}$ (8)
@@
Size of the credal prediction. As discussed above, and further elaborated upon in §E.5.3, a sensible measure of the epistemic uncertainty attached to a random-set prediction $\mathbf{\hat{bel}}$ is the size of the corresponding credal set. Several ways exist of measuring the size of a convex polytope such as a credal set (Sale et al., 2023). Given the upper and lower bounds to the probability assigned to each class $c$ by distributions within the predicted credal set $\hat{Cre}$ (Eq. 4),
$\overline{P}(c)=\max_{P\in\hat{Cre}}P(c),\quad\underline{P}(c)=\min_{P\in\hat{Cre}}P(c),\vspace{-5pt}$ (9)
we propose a simple way to measure the size of the credal set as the difference between the lower and upper bounds (Eq. 9) associated with the most likely class, according to the pignistic prediction. Note that the predicted pignistic estimate $BetP(c)$ falls within the interval $[\underline{P}(c),\overline{P}(c)]$ for each class $c$, not just for the most likely class $\hat{c}$. Its width $\overline{P}(c)-\underline{P}(c)$ indicates the epistemic uncertainty associated with the prediction (§A.3).
@@ 4.1 Implementation
Datasets. Our experiments are performed on multi-class image classification datasets, including MNIST (LeCun & Cortes, 2005), CIFAR-10 (Krizhevsky et al., 2009), Intel Image (Bansal, 2019), CIFAR-100 (Krizhevsky, 2012), and ImageNet (Deng et al., 2009). For out-of-distribution (OoD) experiments, we assess several in-distribution (iD) vs OoD datasets: CIFAR-10 vs SVHN (Netzer et al., 2011)/Intel-Image (Bansal, 2019), MNIST vs F-MNIST (Xiao et al., 2017)/K-MNIST (Clanuwat et al., 2018), and ImageNet vs ImageNet-O (Hendrycks et al., 2021). The data is split into 40000:10000:10000 samples for training, testing, and validation respectively for CIFAR-10 and CIFAR-100, 50000:10000:10000 samples for MNIST, 13934:3000:100 for Intel Image, 1172498:50000:108669 for ImageNet. For OoD datasets, we use 10,000 testing samples, except for Intel Image (3,000) and ImageNet-O (2,000). Training images are resized to $224\times 224$ pixels.
Baselines, backbone and training details. Our baselines include state-of-the-art Bayesian methods LB-BNN (Hobbhahn et al., 2022) and FSVI (Rudner et al., 2022), Ensemble classifiers DE (5 ensembles) (Lakshminarayanan et al., 2017) and ENN (3 ensembles) (Osband et al., 2024), and standard CNN (see Sec. 4.2. Our baseline models are similar to those used in the most recent work (Cinquin & Bamler, 2024; Daxberger et al., 2021; Wu & Williamson, 2024) in uncertainty estimation. All models, including RS-NN, are trained on ResNet50 (on NVIDIA A100 80GB GPUs) with a learning rate scheduler initialized at 1e-3 with 0.1 decrease at epochs 80, 120, 160 and 180. Standard data augmentation (Krizhevsky et al., 2012), including random horizontal/vertical shifts with a magnitude of 0.1 and horizontal flips, is applied to all models.
@@
ResNet50, excluding the top classification layer, serves as the common architecture. ResNet was originally designed for classification on the ImageNet dataset (1000 classes). To accommodate a reduced number of classes as in smaller datasets (e.g., CIFAR-10/MNIST with 10 classes, Intel Image with 7 classes, and CIFAR-100 with 100 classes), two additional dense layers (1024 and 512 neurons, ReLU activation) are added to ResNet50. Similar techniques are commonly applied in deep learning to adapt model architectures to datasets (He et al., 2016; Zagoruyko & Komodakis, 2016). The output layer of RS-NN on ResNet50 has the same number of units as the number of (selected) focal sets $|\mathcal{O}|$, and uses sigmoid activation, since the ground truth encoding resembles a multi-label classification problem (see Sec. 3.2). For all other models, the final output layer simply consists of a softmax activation for multi-class classification. All the models were trained from scratch for 200 epochs (recommended by most), with a batch size of 128. Our objective is to ensure a fair comparison across all models for all experiments. Tuning each model separately to maximise performance would not guarantee that, as some models use pre-trained weights while others train for larger number of epochs, which could result in the over-training of another model. We used pre-trained weights for all models only when trained on ImageNet for efficiency. For all other training hyperparameters (e.g, optimizer), we use what is specified for each model in their respective papers (more in §D, Tab. 5). Consequently, the performance metrics reported in the original papers may vary from those presented here, as they were obtained under different training conditions.
@@
Training the RS-NN. In the clustering phase, we obtain features directly from a pre-trained ResNet50 classifier. We use 50 CPU cores for t-SNE dimensionality reduction, and 150 CPU cores for computing the overlap of classes. We set a budget $K$ of 20 focal sets (ablation study on K in §E.7) for CIFAR-10/ MNIST/ Intel Image, 200 for CIFAR-100 and 3000 for ImageNet. RS-NN is trained from scratch on ground-truth belief encoding of sets using the $\mathcal{L}_{RS}$ loss function (Eq. 7) over 200 epochs, with a batch size ($b_{size}$) of 128 and $\alpha=\beta=1e-3$ as hyperparameter values.
Experiments. Firstly, we assess the test accuracy (%) and inference time (ms) of all baselines on all the multi-class classification datasets. Tab. 1 provides a comparison of test accuracies and inference times of RS-NN against the baselines. Our second set of experiments (Sec. 4.3) concerns out-of-distribution (OoD) detection. Tab. 2 shows our results on OoD detection metrics AUROC (Area Under Receiver Operating Characteristic curve) and AUPRC (Area Under Precision-Recall curve) for all the models on the iD vs OoD datasets listed above. We also report the Expected Calibration Error (ECE) for all models on datasets CIFAR-10, MNIST and ImageNet (see Tab. 2). In a third set of experiments (Sec. 4.4), we test the uncertainty estimation capabilities of RS-NN using both the entropy of the pignistic prediction and the size of the predicted credal set. Tab. 2 presents the predicted pignistic entropy of RS-NN alongside entropies of all baselines on iD and OoD datasets. Tab. 3 shows credal set widths for RS-NN predictions across the same datasets. In our final set of experiments, we explore the scalability of RS-NN (Sec. 4.5) to large-scale architectures, employing it on models such as WideResNet-28-10, VGG16, Inception V3, EfficientNetB2 and ViT-Base-16. Further, to underscore the model’s ability to leverage transfer learning, we train and test on a pre-trained ResNet50 model initialised with ImageNet weights (Tab. 4).
Additional experiments concern obtaining statistical guarantees for RS-NN using non-conformity scores (§A), evaluating the robustness of RS-NN to adversarial attacks (§E.4), noisy and rotated in-distribution samples (Appendix §E.3), showing how RS-NN circumvents the overconfidence problem in CNNs (§E.2). Results showing how credal set width is less correlated with confidence scores than entropy are discussed in in §E.5.3. We also conduct ablation studies on $\alpha$ and $\beta$E.6) and the number of non-singleton focal sets $K$E.7), which show that the best accuracy is obtained at small values ($31e-$ 3) of $\alpha$ and $\beta$, and for $K=20$ (on the CIFAR-10 dataset).
RS-NN (ours) $\mathbf{99.71}\pm\mathbf{0.03}$ $\mathbf{93.53}\pm\mathbf{0.09}$ $\mathbf{94.22}\pm\mathbf{0.03}$ $\mathbf{71.61}\pm\mathbf{0.07}$ $\mathbf{79.92}$ $\mathbf{94.47}$ $\mathbf{1.91}\pm\mathbf{0.02}$
LB-BNN (Hobbhahn et al., 2022) $99.58\pm 0.04$ $89.95\pm 0.81$ $90.49\pm 0.42$ $59.89\pm 1.96$ $72.48$ $90.85$ $7.11\pm 0.89$
FSVI (Rudner et al., 2022) $99.18\pm 0.03$ $80.29\pm 0.05$ $88.92\pm 0.13$ $53.34\pm 0.09$ $62.56$ $84.69$ $340.25\pm 0.76$
DE (Lakshminarayanan et al., 2017) $99.25\pm 0.01$ $92.73\pm 0.04$ $91.98\pm 0.11$ $70.53\pm 0.07$ $78.77$ $94.37$ $13163.50\pm 3.37$
ENN (Osband et al., 2024) $99.07\pm 0.11$ $91.55\pm 0.60$ $91.49\pm 0.19$ $68.02\pm 0.26$ $71.82$ $89.48$ $3.10\pm 0.03$
CNN $99.12\pm 0.04$ $92.08\pm 0.42$ $90.89\pm 0.10$ $65.50\pm 0.08$ $78.56$ $94.34$ $\mathbf{1.91}\pm\mathbf{0.03}$
@@
FSVI $80.29$ Predictive Entropy $0.118\pm 0.563$ $0.0521$ $80.59$ $80.84$ $0.413\pm 0.461$ $74.27$ $72.51$ $0.289\pm 0.670$
DE $92.73$ Mean Entropy $0.154\pm 0.367$ $\mathbf{0.0482}$ $93.84$ $91.88$ $0.939\pm 0.554$ $94.25$ $79.36$ $1.166\pm 0.552$
@@
RS-NN $\mathbf{99.71}$ Pignistic entropy $0.010\pm 0.111$ $\mathbf{0.0029}$ $\mathbf{93.89}$ $\mathbf{93.98}$ $0.530\pm 0.770$ $\mathbf{96.75}$ $\mathbf{96.58}$ $\mathbf{0.740}\pm\mathbf{0.917}$
@@
FSVI $99.18$ Predictive Entropy $0.006\pm 0.265$ $0.0047$ $92.79$ $91.17$ $0.264\pm 0.289$ $91.65$ $95.75$ $0.313\pm 0.381$
DE $99.25$ Mean Entropy $0.031\pm 0.155$ $0.0031$ $92.30$ $92.05$ $\mathbf{0.584}\pm\mathbf{0.587}$ $95.81$ $94.71$ $0.564\pm 0.715$
ENN $99.07$ Mean Entropy $0.022\pm 0.127$ $0.0039$ $81.79$ $82.92$ $0.313\pm 0.464$ $95.94$ $95.45$ $0.503\pm 0.672$